OpenStudy (anonymous):

Gretchen and Ezia received equal scores on a test made up of multiple choice questions and an essay. Gretchen got 18 multiple choice questions correct and received 19 points for her essay. Ezia got 15 multiple choice questions correct and received 31 points for her essay. How many points was each multiple choice question worth?

OpenStudy (anonymous):

you will have to set this up into two equations and then use substitution or elimination. so the first equation would be 18x+19=y can you figure out the second equation from here?

OpenStudy (anonymous):

no see i dont understand it so i'm trying to learn plz help me understand

OpenStudy (anonymous):

ok do you know what substitution and elimination are?

OpenStudy (anonymous):

i think the substitutions are the letters and the elimination is wehn you have to use multiplication or division on both side the goal is t to get x by it self i think i may be wrong

OpenStudy (anonymous):

We will just use substitution for this problem. The method of solving "by substitution" works by solving one of the equations (you choose which one) for one of the variables (you choose which one), and then plugging this back into the other equation, "substituting" for the chosen variable and solving for the other. Then you back-solve for the first variable.

OpenStudy (anonymous):

do you know how I got the equation 18x+19=y

OpenStudy (anonymous):

well i know you got 18 and 19 from the paragraph

OpenStudy (anonymous):

ok thats a start. I choose the numbers that had to do with gretchen. Can you tell me the numbers that have to do with ezia

OpenStudy (anonymous):

Gretchen - 18 and 19 Ezia - 15 and 31

OpenStudy (anonymous):

yep good job now which numbers from both people have to do with multiple choice questions?

OpenStudy (anonymous):

Gretchen -18 Ezia - 15

OpenStudy (anonymous):

good now as you say in the first equation I gave you put the x next to those numbers, each in there own equation. Then you will add the points from the essay question to each of those in there own equation.

OpenStudy (anonymous):

can you show me this?

OpenStudy (anonymous):

18=15

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

not quite. remember the equation I gave you at the very beginning?

OpenStudy (anonymous):

oh yea so 18x+19=y

OpenStudy (anonymous):

but how do i know if they are talking about Gretchen or Ezia

OpenStudy (anonymous):

You know by looking at the question and the sentence starts with their name. for example Gretchen got 18 multiple choice questions correct and received 19 points for her essay.

OpenStudy (anonymous):

well when they asked the question though it didn't give a name "How many points was each multiple choice question worth?"

OpenStudy (anonymous):

that is what we are trying to figure out so that will be our x

OpenStudy (anonymous):

the number u have to divide by

OpenStudy (anonymous):

sorry i didn't explain that well enough. it is in the description that has all of the names and information. and the question tells you what you are trying to figure out and we will put x in for the worth of the points in the equation

OpenStudy (anonymous):

ooooh ok

OpenStudy (anonymous):

so can you set up the second equation using the first one I gave you as a guide?

OpenStudy (anonymous):

both equations will have the same components just different numbers that correspond to that student

OpenStudy (anonymous):

ok man im tired we are going to go back over this in class and i need to finish this assessment thx for your help but whats the answer

OpenStudy (anonymous):

15x +31=y 18x +19=y 18x+19=15x+31 3x=12 x=4 which is your answer

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yep hopefully your teacher can explain better than I could

OpenStudy (anonymous):

@Jennaluv

OpenStudy (anonymous):

ty